Conceptual Example 14 provides useful background for this problem. A playground carousel is free to rotate about its center on frictionless bearings, and air resistance is negligible. The carousel itself (without riders) has a moment of inertia of 102 kg·m2. When one person is standing at a distance of 1.42 m from the center, the carousel has an angular velocity of 0.626 rad/s. However, as this person moves inward to a point located 0.520 m from the center, the angular velocity increases to 0.877 rad/s. What is the person's mass?
Answer #1

Suppose the person's mass = M

So, initial moment of inertia of carousel, I1 = 102 kg*m^2 + M*1.42^2

Initial angular velocity, w1 = 0.626 rad/s

Final moment of inertia, I2 = 102 kg*m^2 + M*0.52^2

Final angular velocity, w2 = 0.877 rad/s

Aply conservation of angular momentum -

I1*w1 = I2*w2

=>(102 + M*1.42^2)*0.626 = (102 + M*0.52^2)*0.877

=> 63.85 + 1.26*M = 89.45 + 0.24*M

=> 1.02*M = 25.6

=> M = 25.6/1.02 = 25.10 kg

So, the person's weight = 25.10 kg (Answer)